Suppose that a recent poll found that 66​% of adults believe that the overall state of moral values is poor. Complete parts​ (a) through​ (c). ​(a) For 350 randomly selected​ adults, compute the mean and standard deviation of the random variable​ X, the number of adults who believe that the overall state of moral values is poor. The mean of X is nothing.​ (Round to the nearest whole number as​ needed.) The standard deviation of X is nothing. ​(Round to the nearest tenth as​ needed.) ​(b) Interpret the mean. Choose the correct answer below. A. For every 231 ​adults, the mean is the maximum number of them that would be expected to believe that the overall state of moral values is poor. B. For every 350 ​adults, the mean is the number of them that would be expected to believe that the overall state of moral values is poor. C. For every 350 ​adults, the mean is the minimum number of them that would be expected to believe that the overall state of moral values is poor. D. For every 350 ​adults, the mean is the range that would be expected to believe that the overall state of moral values is poor. ​(c) Would it be unusual if 230 of the 350 adults surveyed believe that the overall state of moral values is​ poor? No Yes
Answer #1

Sample size , n =    350
Probability of an event of interest, p =   0.66

a)
Mean = np = 350 * 0.66 = 231
std dev ,σ=√np(1-p)=   8.9

b)

For every 350 ​adults, the mean is the number of them that would be expected to believe that the overall state of moral values is poor.

c)

P(X ≤   230   )                  
                          
Z=(X - µ ) / σ = (   230   -   231   ) /   8.8623   =   -0.113
                          
=P(Z≤   -0.113   ) =    0.4551              

No,  it would not be unusual if 230 of the 350 adults surveyed believe that the overall state of moral values is​ poor